A compound that is 31.9% K and 28.9% Cl by mass decomposes when heated to give O2 and a

compound that is 52.4% K and 47.6% Cl by mass. Write the balanced chemical equation for the

reaction.


1
Expert's answer
2020-12-30T07:30:24-0500

In 100 g of stuff, there are 31.9 g potassium, 28.9 g of chlorine, and {[100-(28.9+31.9)]= 39.2 g} oxygen

Moles of potassium = 31.9/39.1= 0.816moles

Moles of chlorine = 28.9/35.5 =

0.816 moles

moles of oxygen = 39.2/16 = 2.45 moles


so empirical formula is KClO3

and reaction is KClO3"\rightarrow" O2 + KCl

a compound that is 52.4% K and 47.6% Cl by mass

It is KCl.